OpenStudy (anonymous):

Can someone explain how to solve this: b ^ 3/2 = 8

OpenStudy (amistre64):

square both sides then cbrt it ....

OpenStudy (anonymous):

b^3/2=8 b=8^2/3 b=2

OpenStudy (anonymous):

sorry b=4

OpenStudy (anonymous):

When having b ^ 2 = 4 where the solution is b = ± 2 why can't we exponentiate both sides to the 1/2 ?

OpenStudy (anonymous):

we can... but this is the easier method

OpenStudy (anonymous):

If you do you ll get b = 4 ^ 1/2 which equals 2 not -2 .. No ?

OpenStudy (amistre64):

quadratics do not have an inverse unless you restrict the domain

OpenStudy (anonymous):

So, why is it correct to do so when having b ^ 3/2 = 8 ?

OpenStudy (amistre64):

you are squaring, squaring is a function such that for any given input there is only 1 result. trying to undo the square is not a function in that for any given input other than 0, there are 2 possible solutions.

OpenStudy (amistre64):

(-2)^2 = 4 (2)^2 = 4 each input has only 1 solution. but we cannot say that sqrt(4) has only 1 solution unless we restrict the domain.

OpenStudy (anonymous):

substitute the values into the equation and you'll get one answer

OpenStudy (anonymous):

Hmm... if understood correctly when having b ^ 3/2 = 8 <=> b ^ 3 = 8 ^ 2. But then why it can be b = 3rd root of 8 ^ 2 and not - 3rd root of 8 ^ 2

OpenStudy (amistre64):

cbrt is a function with an inverse

OpenStudy (amistre64):

there are no 2 input to b^3 that give the same value. therefore we can cbrt any number and get a unique result

OpenStudy (anonymous):

Thnx

OpenStudy (amistre64):

youre welcome